# Convergence order of the quantization error for self-affine measures on Lalley-Gatzouras carpets

### v1

@@ 1.4. Statement of the main results
Unlike the Bedford-McMullen carpets, the following cases are possible:
@@
    for some $\sigma,\omega\in\Psi_{l}$, $|\sigma|\neq|\omega|$, even if $\sigma_{L}=\omega_{L}$; this makes the auxiliary measure in [30]—a Bernoulli product, fail to work. We will construct suitable auxiliary probability measures by applying Prohorov’s theorem.
@@
    for some $\sigma,\omega\in\Psi^{*}$ with $\sigma_{y}\precneqq\omega_{y}$, we have $|\sigma_{L}|>|\omega_{L}|$ (cf. Example 2.1). This makes the method in [20, Lemma 2] no longer applicable, in case $\min_{j\in G_{y}}n_{j}=1$ and $\max_{j\in G_{y}}n_{j}\geq 2$. We will present a completely new method to construct pairwise disjoint approximate squares so that the quantization error for $\mu$ can be estimated from below. For this purpose, we will apply some ideas from [22, p. 695].
@@
Let $(f_{ij})_{(i,j)\in G}$ be as defined in (1.3) and $\mu$ the self-affine measure associated with $(f_{ij})_{(i,j)\in G}$ and a positive probability vector $(p_{ij})_{(i,j)\in G}$. Let $s_{r}$ be the unique positive number satisfying
(1.5) $\lim\limits_{l\to\infty}\frac{1}{l}\log\sum_{\sigma\in\Psi_{l}}(p_{\sigma_{L}}q_{\sigma_{R}}a_{\sigma_{L}}^{r})^{\frac{s_{r}}{s_{r}+r}}=0.$
@@
(1) The existence of the limit in (1.5) and that of the unique number $s_{r}$ will be proved in Section 3. (2) As a consequence of Theorem 1.2, we obtain that, $D_{r}(\mu)=s_{r}$ and the $n$th quantization error for $\mu$ is of the same order as $n^{-\frac{1}{s_{r}}}$. This substantially generalizes our previous work in [20, 30].
The remaining part of the paper is organized as follows. In Section 2, we present some basic facts on the approximate squares which will be used frequently; Using these facts, we establish some characterizations for the quantization error. In section 3, we are devoted to the construction of auxiliary measures by applying Prohorov’s theorem. In section 4, we establish estimates for the quantization coefficient via the auxiliary measures and complete the proof for Theorem 1.2.
@@ 2.1. Some basic facts
Let us begin with an example showing that when $\sigma_{L},\sigma_{\omega}$ are incomparable and $\sigma_{y}\precneqq\omega_{y}$, it can happen that $|\sigma_{L}|>|\sigma_{\omega}|$.
Example 2.1.
Let $m=3,n_{1}=n_{2}=2,n_{3}=1$. Let
$a_{11}=a_{22}=a_{13}=\frac{1}{9};\;a_{21}=a_{12}=\frac{1}{27};\;b_{1}=b_{2}=b_{3}=\frac{1}{3};$
$c_{11}=c_{13}=0;\;c_{21}=\frac{26}{27};\;c_{12}=\frac{1}{9},\;c_{22}=\frac{4}{27};\;d_{1}=0,\;d_{2}=\frac{1}{3};\;d_{3}=\frac{2}{3}.$
For $x\in\mathbb{R}$ and $p\in\mathbb{N}$, we denote by $\{x\}^{p}$ the word of length $p$ with all entries equal to $x$. We define
$\sigma_{L}=((2,2),(1,1),(1,1),\ldots,(1,1))\in G^{12},\;\sigma_{R}=\{3\}^{13};$
$\sigma_{\omega}=((1,2),(2,1),(2,1),\ldots,(2,1))\in G^{9},\;\omega_{R}=\{1\}^{3}\ast\{3\}^{16}.$
Not that $a_{\sigma_{L}}=3^{-24}$ and $a_{\omega_{L}}=3^{-27}$ and that $b_{1}=b_{2}=b_{3}=3^{-1}$. One can see that $|\sigma|=25<|\omega|=28$ and $\sigma_{y}\precneqq\omega_{y}$, while $|\sigma_{L}|=12>9=|\omega_{L}|$.
For two words $\sigma,\omega\in\Phi^{*}$, with $\sigma_{L},\omega_{L}$ comparable, we have
@@
Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})$. Suppose that $\sigma_{L}\preceq\omega_{L}$ and $|\sigma|>|\omega|$. Then for some $p_{1}\geq 0$ and $p_{2}\geq 1$, we have
@@
Remark 2.3.
From Lemma 2.2, one can also see that, for $\sigma,\omega\in\Psi^{*}$, we have, either $F_{\sigma}^{\circ}\cap F_{\omega}^{\circ}=\emptyset$, or $F_{\sigma}\subseteq F_{\omega}$, or $F_{\omega}\subseteq F_{\sigma}$. Indeed, if either $y_{\sigma},y_{\omega}$ are incomparable, or $\sigma_{L},\sigma_{\omega}$ are incomparable, then we clearly have $F_{\sigma}^{\circ}\cap F_{\omega}^{\circ}=\emptyset$; otherwise, by Lemma 2.2, we have $F_{\sigma}\subseteq F_{\omega}$, or $F_{\omega}\subseteq F_{\sigma}$.
For convenience, we write
$\underline{a}:=\min_{(i,j)\in G}a_{ij},\;\overline{a}:=\max_{(i,j)\in G}a_{ij};\;\;\underline{b}:=\min_{j\in G_{y}}b_{j},\;\overline{b}:=\max_{j\in G_{y}}b_{j}.$
We clearly have that $\underline{a}<\overline{b}$, but it is possible that $\overline{a}\geq\underline{b}$.
@@
Let $A_{1}:=[\frac{\log\underline{a}}{\log\overline{b}}]+1$ and $A_{2}:=[\frac{\log\underline{b}}{\log\overline{a}}]+1$. Assume that $\sigma,\omega\in\Psi^{*}$ and $F_{\omega}\subseteq F_{\sigma}$. Then
@@
Suppose that $h>A_{1}$. Then $h-1\geq A_{1}\geq 2$. By the definition of $\Psi_{l}$, we deduce
$\prod_{h=1}^{k+h-1}b_{j_{h}}=\prod_{h=1}^{k}b_{j_{h}}\prod_{h=k+1}^{k+h-1}b_{j_{h}}\leq\prod_{h=1}^{l}a_{i_{h}j_{h}}\overline{b}^{A_{1}}<\prod_{h=1}^{l}a_{i_{h}j_{h}}\underline{a}\leq\prod_{h=1}^{l+1}a_{i_{h}j_{h}}.$
@@
$a_{\sigma_{L}}\prod_{h=l+1}^{l+h}a_{i_{h}j_{h}}\leq a_{\sigma_{L}}\overline{a}^{A_{2}}<\prod_{h=1}^{k-1}b_{j_{h}}\underline{b}\leq\prod_{h=1}^{k}b_{j_{h}}.$
@@
Now we give an example to illustrate Lemma 2.4 (ii).
Example 2.5.
In Example 2.1, we redefine $a_{13}=\frac{1}{12},b_{3}=\frac{1}{10},d_{3}=\frac{9}{10}$, and leave all the other parameters unchanged. Let
$\sigma_{L}=((1,1),(1,1),(1,1),\ldots,(1,1))\in G^{9},\;\sigma_{R}=\{1\}^{9}\ast 3;$
$\omega_{L}=((1,1),(1,1),(1,1),\ldots,(1,1))\in G^{10},\;\omega_{R}=\{1\}^{8}\ast 3.$
Note that $a_{\sigma_{L}}=3^{-18}$ and $a_{\omega_{L}}=3^{-20}$. Note that
$b_{\sigma_{y}^{-}}=b_{\omega_{y}^{-}}=3^{-18}\geq a_{\sigma_{L}},a_{\omega_{L}}>b_{\sigma_{y}}=b_{\omega_{y}}=3^{-18}\cdot\frac{1}{10}.$
It follows that $\sigma\in\Psi_{9},\omega\in\Psi_{10}$ and $F_{\omega}\subsetneq F_{\sigma}$, but $|\sigma|=|\omega|=19$.
We end this subsection with the following observation on the length of $\sigma_{R}$ for a word $\sigma\in\Phi^{*}$. We will need it in the characterization for the quantization error.
@@
Let $A_{3}:=\max\limits_{j\in G_{y}}\max\limits_{1\leq i\leq n_{j}}\frac{a_{ij}}{b_{j}}$ and $A_{4}:=\frac{\log A_{3}}{\log\underline{b}}$. For every $\sigma\in\Psi_{1}$, we have $|\sigma_{R}|\geq 1$; for every $\sigma\in\Psi_{l}$, we have $|\sigma_{R}|\geq A_{4}l$.
@@ 2.2. Characterizations for the quantization error
For every $\sigma\in\Psi_{1}$, we define $\sigma^{\flat}:=\theta$. Next, we assume that $l\geq 2$. For $\sigma\in\Psi_{l}$, we write $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})\in\Psi_{l}$. We have
@@
$\mathcal{E}_{r}(\theta):=1;\;\mathcal{E}_{r}(\sigma):=\mu(F_{\sigma})a_{\sigma_{L}}^{r};\;\overline{\eta}_{r}:=\overline{p}\;\underline{q}^{-1}\overline{a}^{r};\;\underline{\eta}_{r}:=\underline{p}\;\underline{q}^{A_{1}}\underline{a}^{r}.$
@@
(2.1) $\underline{\eta}_{r}\mathcal{E}_{r}(\sigma^{\flat})\leq\mathcal{E}_{r}(\sigma)\leq\overline{\eta}_{r}\mathcal{E}_{r}(\sigma^{\flat})<\mathcal{E}_{r}(\sigma^{\flat}).$
@@
$\Lambda_{n,r}:=\{\sigma\in\Psi^{*}:\mathcal{E}_{r}(\sigma^{\flat})\geq\underline{\eta}_{r}^{n}>\mathcal{E}_{r}(\sigma)\};\;\varphi_{n,r}:={\rm card}(\Lambda_{n,r}).$
@@
(1) For every $n\geq 1$, $\Lambda_{n,r}$ is a finite maximal anti-chain in $\Psi^{*}$. As we did in [20, Lemma 1], it is easy to show that $\varphi_{n,r}\asymp\varphi_{n+1}$. Moreover, for every $\sigma\in\Lambda_{n,r}$, we have, $\underline{\eta}_{r}^{n+1}\leq\mathcal{E}_{r}(\sigma)<\underline{\eta}_{r}^{n}$. (2) Let $l_{n,r}:=\min_{\sigma\in\Lambda_{n,r}}|\sigma_{L}|$. We have, $\underline{\eta}_{r}^{l_{n,r}}<\underline{\eta}_{r}^{n}$, so $l_{n,r}\geq n$.
@@
For $\sigma,\omega\in\Lambda_{n,r}$, if $\sigma_{L},\omega_{L}$ are comparable, then $\sigma_{y},\omega_{y}$ are incomparable, because otherwise, by Lemma 2.2, we would have $F_{\sigma}\subseteq F_{\omega}$, or $F_{\sigma}\subseteq F_{\omega}$, contradicting the fact that $\Lambda_{n,r}$ is a finite anti-chain.
In order to establish estimates for the quantization error, we need to construct a finite anti-chain $F_{n,r}$ out of $\Lambda_{n,r}$, such that $\mathcal{E}_{r}(\tau)\asymp\underline{\eta}_{r}^{n}$ for every $\tau\in F_{n,r}$, and for every pair of distinct words $\tau^{(1)},\tau^{(2)}$ in $F_{n,r}$, the following holds:
In [20, Lemma 2], this was done by selecting, for every $\sigma\in\Lambda_{n,r}$, a word $\sigma^{*}$ with $F_{\sigma^{*}}\subset F_{\sigma}$. Unfortunately, the method in [20] strongly relies on the fact that $a_{ij}\equiv n^{-1}\;((i,j)\in G)$. In fact, that method remains valid under the weaker assumption that $a_{ij}\equiv a_{j}\;(1\leq i\leq n_{j})$ for every $j\in G_{y}$, but it is not applicable to general Lalley-Gatzouras case. We will construct $\mathcal{F}_{n,r}$ in a completely different way. Before we proceed with this construction, let us cope with the extreme case that $n_{j}=1$ for every $j\in G_{y}$, where, $\mathcal{F}_{n,r}$ can be defined in a convenient manner.
@@
Assume that $n_{j}=1$ for every $j\in G_{y}$. Let $\sigma,\omega\in\Psi^{*}$ with $\sigma_{y},\omega_{y}$ comparable. Then we have, either $F_{\omega}\subseteq F_{\sigma}$, or $F_{\sigma}\subseteq F_{\omega}$. In particular, for every pair $\sigma,\omega$ of distinct words in $\Lambda_{n,r}$, we have that $\sigma_{y},\omega_{y}$ are incomparable
@@
By the assumptions that $\sigma_{y}\preceq\omega_{y}$ and $n_{j}=1$ for all $j\in G_{y}$, we know that $\sigma_{L},\omega_{L}$ are comparable. The lemma follows easily from Lemma 2.2. ∎
@@
Let $n>2A_{4}^{-1}A_{2}$. Assume that $n_{j}=1$ for every $j\in G_{y}$. For $\sigma\in\Lambda_{n,r}$, we write $\sigma=((1,j_{1}),\ldots,(l,j_{l}),j_{l+1},\ldots,j_{k})$. From Remark 2.7 (2) and Lemma 2.6, we have, $|\sigma_{R}|>2A_{2}$. By Lemma 2.4 (ii), we may define
$\tau_{\sigma}:=\sigma_{L}\ast((1,j_{l+1}),\ldots,(1,j_{l+1+2A2}))\ast(j_{l+2A_{2}+2},\ldots,j_{k},1,m,\ldots,j_{\overline{k}}).$
@@
Next, we assume that $n_{j_{0}}\geq 2$ for some $j_{0}\in G_{y}$. We will construct $F_{n,r}$ in two steps. For the choice of $\overline{\sigma}$ in the first step, we are inspired by [22, p. 695].
Step 1: For every $\sigma\in\Lambda_{n,r}$, we insert a word $\tau_{0}:=((1,j_{0}),(n_{j_{0}},j_{0}))$ immediately after $\sigma_{L}$. That is, for $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})$, we define
@@
We define $\mathcal{B}_{n,r}:=\{\overline{\sigma}:\sigma\in\Lambda_{n,r}\}$. We will prove that, for all large $n$, for every pair of distinct words $\overline{\sigma},\overline{\omega}\in\mathcal{B}_{n,r}$, we have, either (2.2) holds with $\overline{\sigma},\overline{\omega}$ in place of $\tau^{(1)},\tau^{(2)}$ (cf. Lemma 2.11), or $\overline{\sigma}^{*}_{y},\overline{\omega}^{*}_{y}$ are incomparable (cf. Lemma 2.14).
Step 2: For every $\overline{\sigma}\in\mathcal{B}_{n,r}$, we select a word $\overline{\sigma}^{*}$ such that $F_{\overline{\sigma}^{*}}\subset F_{\overline{\sigma}}$, and for every pair $\overline{\sigma}^{*},\overline{\sigma}^{*}$ of distinct words, (2.2) holds with $\widetilde{\sigma}^{*},\widetilde{\omega}^{*}$ in place of $\tau^{(1)},\tau^{(2)}$. Then we define $\mathcal{F}_{n,r}:=\{\overline{\sigma}^{*}:\overline{\sigma}\in\mathcal{B}^{\flat}_{n,r}\}$.
@@
Let $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma_{L},\omega_{L}$ incomparable and $\overline{\sigma}_{y}\preceq\overline{\omega}_{y}$. We have
$d_{h}(F_{\overline{\sigma}},F_{\overline{\omega}})\geq 2^{-\frac{1}{2}}\underline{a}^{2}\max\{|F_{\overline{\sigma}}|,|F_{\overline{\omega}}|\}.$
@@
As $\sigma_{L},\omega_{L}$ are incomparable, we have, $f_{\sigma_{L}}(E_{0}),f_{\omega_{L}}(E_{0})$ are non-overlapping. Note that $F_{\overline{\sigma}}\subset f_{\sigma_{L}\ast\tau_{0}}(E_{0})$ and $F_{\overline{\omega}}\subset f_{\omega_{L}\ast\tau_{0}}(E_{0})$. Using this and the assumption that $\overline{\sigma}_{y}\preceq\overline{\omega}_{y}$, we deduce
@@
In the following we are going to examine the comparability between $\overline{\sigma}_{y},\overline{\omega}_{y}$, when $\sigma_{L},\omega_{L}$ are comparable. We begin with the simplest cases when $|\omega_{L}|-|\sigma_{L}|\leq 2$. By Lemma 2.6 and Remark 2.7 (2), for every $n\geq 4A_{4}^{-1}$ and every $\tau\in\Lambda_{n,r}$, we have $|\tau_{R}|\geq 4$. We will assume that $n\geq 4A_{4}^{-1}$ in the subsequent Lemma 2.12 and Example 2.13 in order to avoid some trivial cases.
@@
Let $\sigma,\omega\in\Lambda_{n,r}$. Assume that $\sigma_{L}\preceq\omega_{L}$ and $|\omega_{L}|-|\sigma_{L}|\leq 2$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.

### v10

@@ 1.5. Statement of the main results
Unlike Bedford-McMullen carpets, we have two major difficulties which arises from the geometric structure of general Lalley-Gatzouras carpets.
@@
    It is possible that, for some $\sigma,\omega\in\Psi_{l}$, $|\sigma|\neq|\omega|$, even if $\sigma_{L}=\omega_{L}$; for this reason, one can not split the sum $\sum_{\sigma\in\Phi_{l}}(p_{\sigma_{L}}q_{\sigma_{L}}a_{\sigma_{L}}^{r})^{t}$ into two independent factors, one of which is a sum over $(i,j)\in G$ and the other is a sum over $j\in G_{y}$ (cf. (1.5)). This makes the auxiliary measure in [26]—a Bernoulli product, fail to work. We will construct auxiliary probability measures by applying Prohorov’s theorem.
@@
    It may happen that, for some $\sigma,\omega\in\Psi^{*}$ with $\sigma_{y}\precneqq\omega_{y}$, we have $|\sigma_{L}|>|\omega_{L}|$. This makes the method in [13, Lemma 2] no longer applicable, in case $\min_{j\in G_{y}}n_{j}=1$ and $\max_{j\in G_{y}}n_{j}\geq 2$. We will present a new method to construct pairwise disjoint approximate squares so that the quantization error for $\mu$ can be estimated from below.
@@
Let $\mathcal{I}=(f_{ij})_{(i,j)\in G}$ be as defined in (1.3) and $\mu$ the self-affine measure associated with $\mathcal{I}$ and a positive probability vector $(p_{ij})_{(i,j)\in G}$. Let $s_{r}$ be the unique positive number satisfying
(1.6) $\lim_{l\to\infty}\frac{1}{l}\log\sum_{\sigma\in\Psi_{l}}(p_{\sigma_{L}}q_{\sigma_{R}}a_{\sigma_{L}}^{r})^{\frac{s_{r}}{s_{r}+r}}=0.$
@@
(1) The existence of the pressure-like function in (1.6) and that of the unique number $s_{r}$ will be proved in Section 3, by applying a sub-additivity argument as used in thermodynamic formulism (cf. [3]). (2) As a consequence of Theorem 1.2, we obtain that, $D_{r}(\mu)=s_{r}$ and the $n$th quantization error for $\mu$ is of the same order as $n^{-\frac{1}{s_{r}}}$. This substantially generalizes our previous work on the quantization for self-affine measures on Bedford-McMullen carpets.
It was mentioned in [14, Example 1.15], that for every $q\in(0,1)$, one can obtain an expression for the $L^{q}$-spectrum $T(q)$ for $\mu$ by means of the formula for the quantization dimension. As an application of Feng-Wang’s result [4, Theorem 2], we will show that, for every $q\in[0,\infty)$, $T(q)$ can be expressed in terms of approximate squares of the same order. That is,
Proposition 1.4.
Let $\mu$ be the same as in Theorem 1.2. For $q\in[0,\infty)$, the $L^{q}$-spectrum $T(q)$ for $\mu$ is the unique solution of the following equation:
$\lim_{l\to\infty}\frac{1}{l}\log\sum_{\sigma\in\Psi_{l}}(p_{\sigma_{L}}q_{\sigma_{R}})^{q}a_{\sigma_{L}}^{T(q)}=0.$
We will complete the proof of Theorem 1.2 in three steps. First, we will prove some basic facts in Section 2 on the approximate squares. Using these facts, we will construct a set of pairwise disjoint approximate squares $F_{\sigma}$ of similar ”energy” $\mathcal{E}_{r}(\sigma):=p_{\sigma_{L}}q_{\sigma_{R}}a_{\sigma_{L}}^{r}$; then we apply the three-step procedure as described in [12] to establish some useful characterizations for the quantization error. Secondly, we will construct some auxiliary measures by applying Prohorov’s theorem in section 3. Finally, in section 4, we establish estimates for the quantization coefficient via the auxiliary measures and complete the proof for the main results.
@@ 2.1. Some basic facts
One can easily construct examples and see that, when $\sigma_{L},\omega_{L}$ are incomparable and $\sigma_{y}\precneqq\omega_{y}$, it can happen that $|\sigma_{L}|>|\sigma_{\omega}|$. In contrast, for two words $\sigma,\omega\in\Psi^{*}$, with $\sigma_{L},\omega_{L}$ comparable, we have
@@
Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})$. Suppose that $|\sigma|>|\omega|$. Then for some $p_{1}\geq 0$ and $p_{2}\geq 1$, we have
@@
For convenience, we define
$\underline{a}=\min\limits_{(i,j)\in G}a_{ij},\;\overline{a}=\max\limits_{(i,j)\in G}a_{ij};\;\;\underline{b}=\min\limits_{j\in G_{y}}b_{j},\;\overline{b}=\max\limits_{j\in G_{y}}b_{j};$
(2.1) $A_{1}=[\frac{\log\underline{a}}{\log\overline{b}}]+1,\;A_{2}=[\frac{\log\underline{b}}{\log\overline{a}}]+1,\;A_{3}=\max\limits_{j\in G_{y}}\max\limits_{1\leq i\leq n_{j}}\frac{a_{ij}}{b_{j}},\;A_{4}=\frac{\log A_{3}}{\log\underline{b}}.$
These constants will be used to control the length of $\tau_{R}$ in the construction of sub-approximate squares $F_{\tau_{L}\ast\tau_{R}}$.
@@
Assume that $\sigma,\omega\in\Psi^{*}$ and $F_{\omega}\subseteq F_{\sigma}$. Then we have
@@
Suppose that $h>A_{1}(\geq 2)$. By the definition of $\Psi_{l}$, we deduce
$\prod_{p=1}^{k+h-1}b_{j_{p}}=\prod_{p=1}^{k}b_{j_{p}}\prod_{p=k+1}^{k+h-1}b_{j_{p}}\leq\prod_{p=1}^{l}a_{i_{p}j_{p}}\overline{b}^{A_{1}}<\prod_{p=1}^{l}a_{i_{p}j_{p}}\underline{a}\leq\prod_{p=1}^{l+1}a_{i_{p}j_{p}}.$
@@
$a_{\sigma_{L}}\prod_{p=l+1}^{l+h}a_{i_{p}j_{p}}\leq a_{\sigma_{L}}\overline{a}^{A_{2}}<\prod_{p=1}^{k-1}b_{j_{p}}\underline{b}\leq\prod_{p=1}^{k}b_{j_{p}}.$
@@
Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})\in\Psi_{l}$. For $\omega\in\Psi_{l+1}$ with $\sigma\preceq\omega$, by Lemma 2.2, for some $1\leq i\leq n_{j_{l+1}}$ and $p\geq 0$, we have
$\omega=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(i,j_{l+1}),j_{l+2},\ldots,j_{k},\ldots,j_{k+p}).$
We end this subsection with the following observation on the length of $\sigma_{R}$ for a word $\sigma\in\Psi^{*}$. We will need it in the characterization for $e_{\varphi_{n,r},r}(\mu)$.
@@
For $\sigma\in\Psi_{1}$, we have $|\sigma_{R}|\geq 1$; for $\sigma\in\Psi_{l}$, we have $|\sigma_{R}|\geq A_{4}l$.
@@ 2.2. Characterizations for the quantization error
In order to establish estimates for the quantization error, we will construct a finite anti-chain $\mathcal{F}_{n,r}$, such that $\mathcal{E}_{r}(\tau)\asymp\underline{\eta}_{r}^{n}$ for every $\tau\in\mathcal{F}_{n,r}$, and for every pair of distinct words $\tau^{(1)},\tau^{(2)}$ in $\mathcal{F}_{n,r}$, the following holds:
For every $\sigma\in\Psi_{1}$, we define $\sigma^{\flat}:=\theta$. Next, we assume that $l\geq 2$. For $\sigma\in\Psi_{l}$, we write $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})$. We have
@@
(2.3) $\mathcal{E}_{r}(\theta):=1;\;\mathcal{E}_{r}(\sigma):=\mu(F_{\sigma})a_{\sigma_{L}}^{r};\;\overline{\eta}_{r}:=\overline{p}\;\underline{q}^{-1}\overline{a}^{r};\;\underline{\eta}_{r}:=\underline{p}\;\underline{q}^{A_{1}}\underline{a}^{3r}.$
@@
(2.4) $\underline{\eta}_{r}\mathcal{E}_{r}(\sigma^{\flat})<\underline{q}^{A_{1}}\underline{a}^{r}\mathcal{E}_{r}(\sigma^{\flat})\leq\mathcal{E}_{r}(\sigma)\leq\overline{\eta}_{r}\mathcal{E}_{r}(\sigma^{\flat})<\mathcal{E}_{r}(\sigma^{\flat}).$
@@
(2.5) $\Lambda_{n,r}:=\{\sigma\in\Psi^{*}:\mathcal{E}_{r}(\sigma^{\flat})\geq\underline{\eta}_{r}^{n}>\mathcal{E}_{r}(\sigma)\};\;\varphi_{n,r}:={\rm card}(\Lambda_{n,r}).$
@@
(1) For every $n\geq 1$, $\Lambda_{n,r}$ is a finite maximal anti-chain in $\Psi^{*}$. As we did in [13, Lemma 1], it is easy to show that $\varphi_{n,r}\asymp\varphi_{n+1}$. Moreover, for every $\sigma\in\Lambda_{n,r}$, we have, $\underline{\eta}_{r}^{n+1}\leq\mathcal{E}_{r}(\sigma)<\underline{\eta}_{r}^{n}$. (2) Let $l_{n,r}:=\min_{\sigma\in\Lambda_{n,r}}|\sigma_{L}|$. We have, $\underline{\eta}_{r}^{l_{n,r}}<\underline{\eta}_{r}^{n}$, so $l_{n,r}\geq n$.
@@
For $\sigma,\omega\in\Lambda_{n,r}$, if $\sigma_{L},\omega_{L}$ are comparable, then $\sigma_{y},\omega_{y}$ are incomparable, because otherwise, by Lemma 2.1, we would have $F_{\sigma}\subseteq F_{\omega}$, or $F_{\omega}\subseteq F_{\sigma}$, contradicting the fact that $\Lambda_{n,r}$ is a finite anti-chain.
Next, we are going to construct the finite anti-chain $\mathcal{F}_{n,r}$ out of $\Lambda_{n,r}$. First, let us cope with the extreme case that $n_{j}=1$ for every $j\in G_{y}$, where, $\mathcal{F}_{n,r}$ can be defined in a convenient manner.
@@
Assume that $n_{j}=1$ for every $j\in G_{y}$. Let $\sigma,\omega\in\Psi^{*}$ with $\sigma_{y},\omega_{y}$ comparable. Then we have, either $F_{\omega}\subseteq F_{\sigma}$, or $F_{\sigma}\subseteq F_{\omega}$. In particular, for every pair $\sigma,\omega$ of distinct words in $\Lambda_{n,r}$, $\sigma_{y},\omega_{y}$ are incomparable.
@@
By the assumptions that $\sigma_{y},\omega_{y}$ are comparable and $n_{j}=1$ for all $j\in G_{y}$, we know that $\sigma_{L},\omega_{L}$ are comparable. The lemma follows from Lemma 2.1. ∎
@@
Let $n>2A_{4}^{-1}A_{2}$. Assume that $n_{j}=1$ for every $j\in G_{y}$. For $\sigma\in\Lambda_{n,r}$, we write $\sigma=((1,j_{1}),\ldots,(1,j_{l}),j_{l+1},\ldots,j_{k})$. From Remark 2.4 (2) and Lemma 2.3, we have, $|\sigma_{R}|>2A_{2}$. By Lemma 2.2 (ii), we may define
$\tau_{\sigma}:=\sigma_{L}\ast((1,j_{l+1}),\ldots,(1,j_{l+2A2}))\ast(j_{l+2A_{2}+1},\ldots,j_{k},1,m,\ldots,j_{\overline{k}}).$
@@
In the following, we assume that $n_{j_{0}}\geq 2$ for some $j_{0}\in G_{y}$. We will construct $\mathcal{F}_{n,r}$ in two steps. For the choice of $\overline{\sigma}$, we are inspired by [13, Lemma 2] and [16, Lemma 7.2].
Step 1: For $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})\in\Lambda_{n,r}$, we define
@@
We define $\mathcal{B}_{n,r}:=\{\overline{\sigma}:\sigma\in\Lambda_{n,r}\}$. We will prove that, for all large $n$, for every pair of distinct words $\overline{\sigma},\overline{\omega}\in\mathcal{B}_{n,r}$, we have, either of the following holds:
(i) $d_{h}(F_{\overline{\sigma}},F_{\overline{\omega}})\gtrsim\max\{|F_{\overline{\sigma}}|,|F_{\overline{\omega}}|\}$ (Lemma 2.8);
(ii) $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable (Lemma 2.10).
Step 2: For every $\overline{\sigma}\in\mathcal{B}_{n,r}$, we select a word $\overline{\sigma}^{*}$ such that
$d_{v}(F_{\overline{\sigma}^{*}},F_{\overline{\omega}^{*}})\gtrsim\max\{|F_{\overline{\sigma}^{*}}|,|F_{\overline{\sigma}^{*}}|\},$
whenever $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable. We define $\mathcal{F}_{n,r}:=\{\overline{\sigma}^{*}:\overline{\sigma}\in\mathcal{B}_{n,r}\}$.
@@
Let $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma_{L},\omega_{L}$ incomparable and $\overline{\sigma}_{y},\overline{\omega}_{y}$ comparable. We have, $d_{h}(F_{\overline{\sigma}},F_{\overline{\omega}})\geq 2^{-\frac{1}{2}}\underline{a}^{2}\max\{|F_{\overline{\sigma}}|,|F_{\overline{\omega}}|\}$.
@@
Let $\tau_{0}=((1,j_{0}),(n_{j_{0}},j_{0}))$. Without loss of generality, we assume that $|\sigma_{L}|=l_{1},|\omega_{L}|=l_{2}$ and that $l_{1}\leq l_{2}$. Since $\overline{\sigma}_{y},\overline{\omega}_{y}$ are comparable, we have that $(\omega_{L})_{y}|_{l_{1}}=(\sigma_{L})_{y}$. As $\sigma_{L},\omega_{L}$ are incomparable, we know that $\omega_{L}|_{l_{1}}\neq\sigma_{L}$. Therefore, $f_{\sigma_{L}}(E_{0}),f_{\omega_{L}|_{l_{1}}}(E_{0})$ are non-overlapping. Note that $F_{\overline{\sigma}}\subset f_{\sigma_{L}\ast\tau_{0}}(E_{0})$ and $F_{\overline{\omega}}\subset f_{\omega_{L}\ast\tau_{0}}(E_{0})\subset f_{\omega_{L}|_{l_{1}}}(E_{0})$. We deduce
@@
In the following, we are going to show that, for $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma_{L}\preceq\omega_{L}$, $\overline{\sigma}_{y},\overline{\omega}_{y}$, are necessarily incomparable. We begin with the simplest cases when $|\omega_{L}|-|\sigma_{L}|\leq 2$. By Lemma 2.3 and Remark 2.4 (2), for every $n\geq 4A_{4}^{-1}$ and every $\tau\in\Lambda_{n,r}$, we have $|\tau_{R}|\geq 4$. We assume that $n\geq 4A_{4}^{-1}$ in the subsequent Lemma 2.9 in order to avoid some trivial cases.
@@
Let $\sigma,\omega\in\Lambda_{n,r}$ be distinct words. Assume that $\sigma_{L}\preceq\omega_{L}$ and $|\omega_{L}|-|\sigma_{L}|\leq 2$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.